-2x+2y=2
-4y-2x=-22
Solve the system of equations

Answer by ewatrrr(24785)

 
Hi
Solve the system of equations. Will use Elimination in this case
Note: -2x+2y=2 is the same as 2y-2x = 2
2y-2x =2 |multiplying thru by -1 to eliminate the x variable
-4y-2x=-22
-2y+2x = -2 |result of multiplying thru by -1 to eliminate the x variable
-4y-2x =-22
-6y = -24
y = 4 and x = 3 (8 = 2x+2, 6 = 2x and x=3)
CHECKING our Answer***Using -4y-2x=-22
-8 -6 = -22